# PTM-Psi on the Cloud

### v1

@@ Abstract
We developed an advanced computational architecture to accelerate Post-Translational Modifications on protein structures and interactions (PTM-Psi) simulations utilizing asynchronous, loosely coupled workflows on the Azure Quantum Elements platform. The cloud architecture harnessed inherent task parallelism to enhance simulation throughput and dynamically allocate computational resources to optimize efficiency. Here, we seamlessly integrate emerging cloud computing assets that further expand the scope and capability of PTM-Psi simulations beyond current limitations. By refactoring the existing PTM-Psi workflow using well-established software, including Alphafold2, NWChem, GROMACS, and Python-based analysis tools into a cloud-native library, we optimized resource allocation tailored to each workflow’s needs for scaling up the PTM-Psi simulation. We employed a “flow-of-workflow” approach in which the cloud architecture supports the computational investigation of a combinatorial explosion of thiol PTMs on an exemplary protein megacomplex critical to the Calvin-Benson cycle of light-dependent sugar production in cyanobacteria. With PTM-Psi on the cloud, we transformed the pipeline for the thiol PTM analysis to achieve high throughput by leveraging the strengths of the cloud service. PTM-Psi on the cloud reduces operational complexity and lowers entry barriers to data interpretation with structural modeling for a redox proteomics mass spectrometry specialist.
@@ 1 Introduction
Since its inception, thiol-based redox proteomics with mass spectrometry (MS), where all protein thiols are subject to a variety of PTMs, has become the primary technology for broad identification and quantification of site-specific redox PTMs to understand the redox regulation in signaling and metabolism under physiological or pathological conditions2. Profiling thiol redox PTMs and understanding their impact on biological functions is an incredibly challenging task due to the combinatorial explosion of possible outcomes. To meet the technical need of data analysis and interpretation with structural models, we developed a user-friendly Python-based package, Post-translational Modifications on Protein Structure, and their impacts on dynamics and functions (PTM-Psi), that parses the complex workflow for numerous combinations of PTMs on proteins. PTM-Psi integrates several open-source software packages into its workflow, including software to infer protein structure from sequence (AlphaFold25), force field development for non-standard amino acids (including redox PTM forms) using quantum chemistry (NWChem6), free energy perturbation methods such as thermodynamic integration through molecular dynamics (GROMACS7), and bound complex docking algorithms (AutoDock8). In addition, the package utilizes a visual tool from ChimeraX9, aiming for an accurate interpretation of experimental data from thiol-based redox proteomics.10
@@
We utilized PTM-Psi to modify cysteines of the homotetramer of glyceraldehyde-3-phosphate dehydrogenase (NADP-GADPH, GAPDH, or GAP2) from Synechococcus elongatus PCC 7942, as an example to interpret a sequence-structure-function relationship derived from thiol redox proteome data10. More specifically, we introduced the S-nitrosothiol (SNO) group through a modification that can be named “S-nitrosylation” or “S-nitrosation”, depending on the specific reaction mechanism and species involved11. We found that, while the impact of SNO moieties on the overall GAPDH conformation remains minimal, introducing an SNO group near the protein surface at cysteine 78 (CYS78), might allosterically inhibit the dehydrogenation reaction at the embedded catalytic site, including cysteine 155 (CYS155), by reducing the affinity towards the cofactor NADP+ due to steric repulsion between the SNO group and the adenosine 2’-phosphate, and the removal of a hydrogen bond between the free thiol of CYS78 and the same adenosine 2’-phosphate.10. We were excited by the new insights from chemistry, modeling, and simulations brought by PTM-Psi in protein sciences. We were eager to further our knowledge about the role that thiol PTMs play in the formation of the protein megacomplex. However, the constraints imposed by the computational resources of standard on-site clusters hindered our ability to scale up our protocol to obtain a predictive understanding of the regulatory mechanisms governed by redox PTMs.
@@
The GAPDH/CP12/PRK dark complex (Figure 1) is an important molecular redox switch of the Calvin-Benson cycle (CBC) in photosynthetic species–such as cyanobacteria, algae, and land plants–as the enzymatic activities of GAPDH and PRK are strongly inhibited when they form part of the ternary supramolecular GAPDH/CP12/PRK complex. In an oxidized state, the two C-terminal cysteines of CP12 form an intramolecular disulfide bridge that forces the intrinsically disordered protein into a hairpin conformation. The partially-folded C-terminal region of CP12 then binds to GAPDH in an obligatory intermediate step with a (GAPDH)-(CP12)2 stoichiometry. An important prerequisite for the GAPDH/CP12 binary complex formation is the presence of NAD+ (not NADP+) inside the binding pocket of GAPDH. Two GAPDH/CP12 binary complexes then bind two oxidized PRK dimers via CP12’s N-terminus, forming the dark complex spontaneously with a final (GAPDH)2-(CP12)4-(PRK)2 stoichiometry 12. Although certain microorganisms possess a CP12 peptide sequence that allows the formation of an additional intramolecular disulfide bond near the N-terminus, the presence of this second disulfide bridge does not appear to be essential for the dark complex formation and stability. In a highly reduced state, the small scaffold protein CP12 becomes intrinsically disordered as a monomeric unit, thus impeding the formation of the supramolecular dark complex. Furthermore, the GAPDH/CP12/PRK complex dissociates in highly reducing environments in a sequence of steps that involve the reduction of the PRK disulfide bonds, the subsequent dissociation of PRK from the ternary complex, the reduction of the CP12 C-terminus disulfide bonds, and the final dissociation of the GAPDH/CP12 binary complex. In their free form, GAPDH and PRK resume enzymatic activity. The regulation mechanism exerted by reversible thiol PTMs on the formation and dissociation of the dark complex remains unclear. This regulation tunes the carbon flow between the photosynthetic reduction cycle in CBC and the oxidative pentose phosphate cycle depending on the light intensity during the natural diel cycle 13.
@@
To understand the fundamental mechanisms of the enzymatic activities in the CBC turnover, the redox proteome of S. elongatus was measured by high-throughput redox mass spectrometry after light disturbance14. The first version of PTM-Psi workflow was developed to provide mass spectrometry specialists with a structural interpretation of their datasets from computational models. However, this workflow could not efficiently handle thousands of simulations stemming from a combinatorial way to modify all thiol groups that might participate in the day/night redox homeostasis of cyanobacteria15. In response to the immense need for high-performance computing, we deployed PTM-Psi on the cloud with the approach of “a flow of workflows” on Microsoft’s Azure Quantum Elements (AQE) cloud computing infrastructure, which we detail in Figure 2A. To manage accompanying data output from these simulations, we also augmented the workflow with a tool for post-processing analyses of massive datasets such as principal component analysis (PCA), dimensionality reduction, clustering, and kinetic transition network (KTN). We introduced a local vs. global scoring function to evaluate the PTM influences on structures and dynamics at a quick glance (Figure 2B).
@@
We demonstrated the performance of PTM-Psi and its new functionality in a computational investigation of several strategically PTMed GAPDH/CP12/PRK dark complex systems from an experimentally resolved mega structure (PDB: 6GVE) of a related cyanobacteria Thermosynechococcus vestitus BP-1. By leveraging the unique advantages of cloud computing, we were able to obtain all simulation instances at a significantly faster rate than other high-performance computing platforms that suffer from heavy usage and prolonged turnaround times. The cysteines (equivalent to CYS78 in the previous study) of GAPDH act like a linchpin that holds the megacomplex together. Their oxidation with sizable thiol groups pivots to destabilize the assembly of the dark complex, even without a significant change in the structure of individual subunits. Our work demonstrates a scalable research capability of PTM-Psi on the cloud that enables a mechanistic understanding of protein sequence, structure, and dynamics related to redox homeostasis at a proteome level.
Figure 1: A structural representation of GAPDH/CP12/PRK megastructure, so-called the “dark complex” (PDB: 6GVE) (A), and a schematic representation of the interaction among the three modules (B). Gray represents GAPDH (or GAP2) tetramers, orange represents CP12, magenta represents PRK dimers, blue represents the NAD cofactor, yellow represents the cysteine targets for PTMs in this study. The capitalized letters in (B) are chain labels in the PDB file, and the line thickness represents the prevalence of the interaction in the simulations of 25 $\mu$s.
@@
Figure 2: (A)Overview of “a flow of workflow” for standing up PTM-Psi on the Cloud. 1) The user can interact through the Azure Quantum Elements Open OnDemand™ interface or the command line. 2) Once a protein target is identified, PTM-Psi either infers a three-dimensional structure by using AlphaFold2 or reads a PDB file. Then, the user can use PTM-Psi to generate hundreds to thousands of combinatorially PTMed structures. 3) These structures are then stored in a hierarchical file with unique identifiers. 4) The template factory then generates all scripts needed to follow the selected workflow. It interfaces with the appropriate hardware and software, including a queueing system either SLURM or PBS. 5) All software called during the workflows is containerized using Apptainer images and automatically pulled from the container registry. 6) The user submits a top-level script to start the workflow, potentially distributed over several nodes worldwide. 7) The workflow is set up to take advantage of the best hardware available for the given task, taking into account the constraints of the software stack. 8) Depending on the selected workflow, PTM-Psi  generates appropriate scripts that perform analyses on the trajectories from the production runs. It includes thermodynamic integration and principal component analysis. (B) The workflow of PTM-Psi10 with the addition of a module for trajectory analysis.
@@ 2.1 Cloud-based HPC cluster
The simulations described here were carried out on Azure Quantum Elements (AQE or Elements) housed and developed by Microsoft Corporation, USA. The AQE platform offers a hybrid computing environment with a wide variety of hardware for high-performance computing (HPC). Each cluster includes utility nodes for scripting, analysis of results, rendering, and visualization through the Open OnDemand™ web-based client16. Depending on the workflow characteristics, the compute nodes can be interconnected with fast fabrics such as InfiniBand. For this research, we utilized two types of nodes: 1) GPU-nodes containing 96 Intel Xeon (Sapphire Rapids) processor cores and 8 H100 GPUs, connected with high-speed NVLink GPU interconnects, per node; and 2) CPU-only nodes containing 120 AMD EPYC 7V12 processor cores per node. The heterogeneous nature of the hardware is suitable for diverse workloads and workflows. Jobs from other traditional HPC platforms are easy to migrate to the AQE through a Slurm queuing system 17, 18 for cloud-based high-performance computing and artificial intelligence (AI).
2.2 PTM-Psi on the cloud and Sample Preparation
We implemented the approach of “a workflow of workflows” when we migrated PTM-Psi to the cloud-based HPC resources (Figure 2 A). By refactoring the existing workflow into a cloud-native library, we optimized resource allocation tailored to the computational needs of each workflow for scaling up the simulations, while establishing an accessible user interface. PTM-Psi is a Python package to facilitate the computational investigation of post-translational modifications on protein structures and their impacts on dynamics and functions in a workflow that uses well-established software, including Alphafold2, NWChem, GROMACS, and python-based analysis tools. (Figure 2 B). We included an additional module of trajectory analysis from the massive datasets.
@@ 2.2 PTM-Psi on the cloud and Sample Preparation
We computationally investigated the impact of selected thiol PTMs on the GAPDH/CP12/PRK complex (“dark complex”) with PTM-Psi. We used the experimentally determined protein megacomplex from Thermosynechococcus vestitus BP-119 (PDBID: 6GVE) . We set the native complex from 6GVE as Ref 0 in Table 1. The NAD cofactor is included in each monomeric GAP2, as found in the 6GVE structure. Ref 0 contains two disulfide bridges in PRK (CYX13-CYS35 and CYX224-CYS230) and two disulfide bridges in CP12 (CYX61-CYX70 and CYS19-CYS29). CYS77 (equivalent to CYS78 in our prior study) in GAP2 is a free cysteine. We systematically oxidized cysteine 77 of GAP2 with S-nitrosylation (SNO), S-sulfhydration (SSH), and S-glutathionylation (SSG) from Ref 1 to 3. We reduced the two pairs of disulfide bridges in PRK from Ref 4 to 9 with SNO, SSH, and SSG. We did the same for CP12 from 10 to 15 in Table 1.
@@
For each instance, the system is solvated with explicit water molecules in a truncated dodecahedron box with a 1 nm padding for all-atom MD simulations, leading to approximately 800k atoms per instance. Enough Na+ and Cl- ions are added to the simulation box to neutralize the net charge and achieve a 0.154 M salt concentration. Minimization then proceeds in a stepwise fashion by restraining the positions of the atoms in the protein complex with decreasing force constants. System equilibration also proceeds in a stepwise fashion. First, the positions of the protein complex are restrained while the solvent is equilibrated at 300 K for 500 ps in the canonical (NVT) ensemble using the stochastic velocity rescaling (V-rescale) thermostat 20, followed by equilibration at 300 K and 1 atm for 500 ps in the isothermal-isobaric (NPT) ensemble with the stochastic cell rescaling (C-rescale) barostat21 in addition to the V-rescale thermostat. Then, the position restraints on the protein complex are lifted, and the NVT and NPT equilibrations are repeated. This ensures a smooth equilibration of the perturbed complex containing the redox PTMs. PTM-Psi was set up to run production NPT molecular dynamics simulations at 300 K and 1 atm for 100 ns using the V-rescale thermostat and C-rescale barostat. All MD simulations constrained the bonds to hydrogen atoms using the LINCS algorithm22 and used a time step of 2 fs.
Ref #                Site (Residue) PTM type

### v2

@@ Abstract
We developed an advanced computational framework to accelerate the study of the impact of post-translational modifications on protein structures and interactions (PTM-Psi) using asynchronous, loosely coupled workflows on the Azure Quantum Elements Cloud platform. We seamlessly integrate emerging cloud computing assets that further expand the scope and capability of PTM-Psi Python package by refactoring it into a cloud compatible library. We employed a “workflow of workflows” approach wherein a parent workflow spawns one or more child workflows, managing them, and acting on their results. This approach enabled us to optimize resource allocation according to each workflow’s needs, and allowed us to use the cloud heterogeneous architecture for the computational investigation of a combinatorial explosion of thiol protein PTMs on an exemplary protein megacomplex critical to the Calvin-Benson cycle of light-dependent sugar production in cyanobacteria. With PTM-Psi on the cloud, we transformed the pipeline for the thiol PTM analysis to achieve high throughput by leveraging the strengths of the cloud service. PTM-Psi on the cloud reduces operational complexity and lowers entry barriers to data interpretation with structural modeling for a redox proteomics mass spectrometry specialist.
@@ 1 Introduction
Since its inception, thiol-based redox proteomics with mass spectrometry, in which all protein thiols are subject to a variety of PTMs, has become the primary technology for broad identification and quantification of site-specific redox PTMs to understand the redox regulation in signaling and metabolism under physiological or pathological conditions2. Profiling thiol redox PTMs and understanding their impact on biological functions is an extremely challenging task because of the combinatorial explosion of possible outcomes. Recent developments in deep learning have advanced the annotation and prediction of PTMs on proteins.5, 6 However, the impact of PTMs on protein function and dynamics remains unclear without mechanistic insights from a structural model. To meet the technical needs of data analysis and interpretation with structural models, we developed a user-friendly Python-based package, Post-translational Modifications on Protein Structure, and their Impacts on dynamics and functions (PTM-Psi), that parses the complex workflow for numerous combinations of PTMs on proteins. PTM-Psi integrates several open-source software packages into its workflows, including software to infer protein structure from sequence (AlphaFold27), force field development for non-standard amino acids (including redox PTM forms) using quantum chemistry (NWChem8), free energy methods such as thermodynamic integration through molecular dynamics (GROMACS9), and bound complex docking algorithms (AutoDock10). In addition, the package uses a visual tool from ChimeraX11, aiming for an accurate interpretation of experimental data from thiol-based redox proteomics.12
@@
In our previous work, we used PTM-Psi to modify cysteines of the homotetramer of glyceraldehyde-3-phosphate dehydrogenase (NADP-GADPH, GAPDH, or GAP2) from Synechococcus elongatus PCC 7942, as an example to interpret a sequence-structure-function relationship derived from thiol redox proteome data12. More specifically, we introduced the S-nitrosothiol (SNO) group through a modification that can be named S-nitrosylation” or S-nitrosation”, depending on the specific reaction mechanism and species involved13. It was observed that, despite the minimal impact of SNO moieties on the global GAPDH conformation, the presence of an SNO group near the protein surface at cysteine 78 (CYS78) could allosterically inhibit the dehydrogenation reaction occurring at the internal catalytic site, specifically including cysteine 155 (CYS155). This inhibition mechanism is predicated on a reduction in the affinity for the cofactor NADP+, attributed to steric repulsion between the SNO group and the adenosine 2’-phosphate, as well as the elimination of a hydrogen bond between the free thiol of CYS78 and the same adenosine 2’-phosphate.12. We were excited by the new insights from chemistry, modeling, and simulations brought by PTM-Psi in protein sciences, including oligomeric regulations.14 We were eager to further our knowledge about the role that thiol PTMs play in the formation and stability of the protein megacomplex composed of GAPDH, chloroplast protein 12 (CP12), and phosphoribulokinase (PRK). However, the constraints imposed by the computational resources of standard on-site clusters hindered our ability to scale up our protocol to obtain a predictive understanding of the regulatory mechanisms governed by redox PTMs in an expedited timeframe.
@@
The GAPDH/CP12/PRK dark complex (Figure 1) is an important molecular redox switch of the Calvin-Benson cycle (CBC) in photosynthetic species–such as cyanobacteria, algae, and land plants–as the enzymatic activities of GAPDH and PRK are strongly inhibited when they form part of the ternary supramolecular GAPDH/CP12/PRK complex. In an oxidized state, the two C-terminal cysteines of CP12 form an intramolecular disulfide bridge that forces the intrinsically disordered protein into a hairpin conformation. The partially-folded C-terminal region of CP12 then binds to GAPDH in an obligatory intermediate step with a (GAPDH)-(CP12)2 stoichiometry. An important prerequisite for the GAPDH/CP12 binary complex formation is the presence of NAD+ (not NADP+) inside the binding pocket of GAPDH. Two GAPDH/CP12 binary complexes then bind two oxidized PRK dimers via CP12’s N-terminus, forming the dark complex spontaneously with a final (GAPDH)2-(CP12)4-(PRK)2 stoichiometry 15. Although certain microorganisms possess a CP12 peptide sequence that allows the formation of an additional intramolecular disulfide bond near the N-terminus, the presence of this second disulfide bridge does not appear to be essential for the dark complex formation and stability. In a highly reduced state, the small scaffold protein CP12 becomes intrinsically disordered as a monomeric unit, thus impeding the formation of the supramolecular dark complex. Furthermore, the GAPDH/CP12/PRK complex dissociates in highly reducing environments in a sequence of steps that involve the reduction of the PRK disulfide bonds, the subsequent dissociation of PRK from the ternary complex, the reduction of the CP12 C-terminus disulfide bonds, and the final dissociation of the GAPDH/CP12 binary complex. In their free form, GAPDH and PRK resume enzymatic activity. The regulation mechanism exerted by reversible thiol PTMs on the formation and dissociation of the dark complex remains unclear. This regulation tunes the carbon flow between the photosynthetic reduction cycle in CBC and the oxidative pentose phosphate cycle, depending on the light intensity during the natural diel cycle 16.
@@
To understand the fundamental mechanisms of the enzymatic activities in the CBC turnover, the redox proteome of S. elongatus was measured by high-throughput redox mass spectrometry after light disturbance17. The first version of PTM-Psi was developed to provide mass spectrometry specialists with a structural interpretation of their datasets from computational models. However, this version could not efficiently orchestrate thousands of simulations stemming from a combinatorial way to modify all thiol groups that might participate in the day/night redox homeostasis of cyanobacteria18. In response to the immense need for high-performance computing (HPC), we deployed PTM-Psi on the cloud with the approach of “workflow of workflows” on a private preview version of Microsoft’s Azure Quantum Elements (AQE) cloud computing infrastructure, which we visually describe in Figure 2A. To manage accompanying data output from these simulations, we also augmented the package with a tool for post-processing analyses of massive datasets that comprises methods such as principal component analysis (PCA), dimensionality reduction, clustering, and kinetic transition network (KTN). We also introduced a local vs. global scoring function to evaluate the PTM influence on structures and dynamics at a quick glance (Figure 2B).
@@
We demonstrated the performance of PTM-Psi and its new functionality in a computational investigation of several strategically PTMed GAPDH/CP12/PRK dark complex systems from an experimentally resolved mega structure (PDB: 6GVE) of the related cyanobacteria Thermosynechococcus vestitus BP-1 (NCBI:txid197221). By leveraging the unique advantages of cloud computing, we were able to obtain all simulation instances at a significantly faster rate than other HPC platforms that suffer from heavy usage and prolonged turnaround times. The cysteines (equivalent to CYS78 in the previous study) of GAPDH act like a linchpin that holds the megacomplex together. Their oxidation with sizable thiol groups pivots to destabilize the assembly of the dark complex, even without a significant change in the structure of individual subunits. Our work demonstrates a scalable research capability of PTM-Psi on the cloud that enables a mechanistic understanding of protein sequence, structure, and dynamics related to redox homeostasis at a proteome level.
[…]
@@
Figure 2: (A)Overview of “a flow of workflow” for standing up PTM-Psi on the cloud. 1) The user can interact through the Azure Quantum Elements Open OnDemand™ interface or the command line. 2) Once a protein target is identified, PTM-Psi either infers a three-dimensional structure by using AlphaFold2 or reads a PDB file. Then, the user can use PTM-Psi to generate hundreds to thousands of combinatorially PTMed structures. 3) These structures are then stored in a hierarchical file with unique identifiers. 4) The template factory then generates all scripts needed to follow the selected workflow. It interfaces with the appropriate hardware and software, including a queueing system, either SLURM or PBS. 5) All software called during the workflows is containerized using Apptainer images and automatically pulled from the container registry. 6) The user submits a top-level script to start the workflow, potentially distributed over several nodes worldwide. 7) The workflow is set up to take advantage of the best hardware available for the given task, considering the constraints of the software stack. 8) Depending on the selected workflow, PTM-Psi generates appropriate scripts that perform analyses on the trajectories from the production runs. It includes thermodynamic integration and PCA. (B) The workflow of PTM-Psi with the addition of a module for trajectory analysis.
@@ 2.1 Cloud-based HPC cluster
The simulations described here were carried out on the AQE platform, housed and developed by Microsoft Corporation, USA. AQE offers a hybrid computing environment with a wide variety of hardware for HPC through Azure Virtual Machines (VMs). An AQE cluster can be configured with several types of VMs, each optimized for a specific purpose. Furthermore, each cluster includes utility nodes for scripting, analysis of results, rendering, and visualization through the Open OnDemand™ web-based client19. Depending on the workflow characteristics, the compute nodes can be interconnected with fast fabrics such as InfiniBand or NVLink. For this research, we used two types of nodes: 1) graphics processing unit (GPU)-nodes containing 96 Intel Xeon (Sapphire Rapids) processor cores and eight H100 GPUs, connected with high-speed NVLink GPU interconnects, per node (ND_H100_v5 series); and 2) CPU-only nodes containing 120 AMD EPYC 7V12 processor cores per node, 4 GB of RAM per core, and a 200 Gb/s Mellanox HDR InfiBand interconnect (HBv2 series). PTM-Psi was also tested on other Azure VM families, including the NCv3 series featuring V100 GPUs and the NC_A100_v4 series with A100 GPUs. The heterogeneous nature of the Cloud hardware architecture is suitable for diverse workloads and workflows. Furthermore, jobs from other traditional HPC platforms are easy to migrate to the AQE through a SLURM queuing system 20, 21.
2.2 Sample preparation
@@
We computationally investigated the impact of selected thiol PTMs on the GAPDH/CP12/PRK complex (“dark complex”) with PTM-Psi. We used the experimentally determined protein megacomplex from Thermosynechococcus vestitus BP-122 (PDBID: 6GVE). We set the native complex from 6GVE as Simulation (Sim.)0 in Table 1. The NAD cofactor is included in each monomeric GAPDH, as found in the 6GVE structure. Sim.0 contains two disulfide bridges in PRK (CYX13-CYX35 and CYX224-CYS230) and two disulfide bridges in CP12 (CYX61-CYX70 and CYS19-CYS29). We note that CYS77 (equivalent to CYS78 in our prior study) in GAPDH is a free cysteine. We systematically oxidized cysteine 77 of GAPDH with S-nitrosylation (SNO), S-sulfhydration (SSH), and S-glutathionylation (SSG) in Sim.1 to 3. We broke the two pairs of disulfide bridges in PRK in Sim.4 to 9 by introducing SNO, SSH, and SSG modifications in both cysteines involved in the original disulfide bridge. We did the same for CP12 in Sim. 10 to 15 in Table 1.
@@
For each simulation instance, the system is solvated with explicit water molecules in a truncated dodecahedron box with a 1 nm padding for all-atom molecular dynamics (MD) simulations, leading to approximately 800k atoms per instance. Enough Na+ and Cl- ions are added to the simulation box to neutralize the net charge and achieve a 0.154 M salt concentration. Minimization then proceeds in a stepwise fashion by restraining the positions of the atoms in the protein complex with decreasing force constants. System equilibration also proceeds in a stepwise fashion. First, the positions of the protein complex are restrained while the solvent is equilibrated at 300 K for 500 ps in the canonical (NVT) ensemble using the stochastic velocity rescaling (V-rescale) thermostat 23, followed by equilibration at 300 K and 1 atm for 500 ps in the isothermal-isobaric (NPT) ensemble with the stochastic cell rescaling (C-rescale) barostat24 in addition to the V-rescale thermostat. Then, the position restraints on the protein complex are lifted, and the NVT and NPT equilibrations are repeated. This ensures a smooth equilibration of the perturbed complex containing the redox PTMs. PTM-Psi was set up to run production NPT molecular dynamics simulations at 300 K and 1 atm for 100 ns using the V-rescale thermostat and C-rescale barostat. All MD simulations constrained the bonds to hydrogen atoms using the LINCS algorithm25 and used a time step of 2 fs.
Sim.#        Site (Residue) PTM type